• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Reitz

    Log entry time 00:04:24 on October27,2001

    Entry number 75107

    keyword=swing shift summary

    16:00 no beam
    17:30 ended run 1238, changed some delays for rich, downloaded trigger, changed to
    withrich configuration and started another cosmic run
    19:15 mcc restored beam, we put BeO in and asked for 5 uA
    unrastered to check beam position, was 0/0 at BPMA and
    19:36 start second part of the waterfall current scan

    RunID Current NoEvents Comment
    1240 77uA 2M PS 1/1/1/1/1
    1242 77uA 1.8M 400/20/2000/200/1
    1243 100uA 1.3M 1/1/1/1/1
    1244 100uA 2.2M 500/30/2400/2500/1

    20:25 MCC tried to reach 120 uA, but experienced a lot of
    beam trips, and decided, that they are only able to deliver 80-uA
    => We decided to go to the white spectrum scan for HRS right
    20:55 during cycling of right Q2, it tripped, tech on call came in to
    reset it, therefore we needed an access
    22:40 magnets are fine, beam is back, starting white spectra scan on right arm

    runid noevents po_left po_right target comments
    1245 1.5M 4120 1114 water PS=1 run
    1246 2.0M 4120 1114 water normal run
    1247 1.0M 4120 1114 water normal run

    changed momenta to 4.12 / 1.098 Gev

    1248 1.0M 4120 1098 water PS=1 run
    1249 1.0M 4120 1098 water normal run

    next shift continues